A rancher is planning to put up 220 yards of fencing. In the morning she puts up 80 yards, and in the afternoon she puts up 40% of the remaining fence. What percent of the fence did she put up that day?
Answer by edjones(7311)

220-80=140
140*.4=56
140+56=196
196/220=.8909
=89.09%
.
Ed